OpenStudy (anonymous):

Which of the following shows a number as a product of prime numbers in exponential form? Choose all that apply. Select one or more: a. 32 · 5 · 7 b. 42 · 5 · 7 c. 33 · 8 d. 33 · 11

OpenStudy (anonymous):

@bibby

OpenStudy (phi):

33 is 3*11 so c and d don't seem correct

OpenStudy (phi):

42 is 6*7 = 2*3*7 that is not promising. so b is out.

OpenStudy (phi):

32 is 2*16 = 2*2*8 and if we keep simplifying we get to 32= 2*2*2*2*2= \( 2^5\) that is a prime number with an exponent, so I would guess that is what they want so choice a: \( 2^5 \cdot 5 \cdot 7 \)

OpenStudy (phi):

Though normally you would write the prime and its exponent, and not disguise it as e.g. 32

OpenStudy (anonymous):

thank you:) @phi